OpenStudy (anonymous):

I need some help checking & evaluating this: \[\huge\int\limits_{0}^{\infty}\frac{e^x}{e^{2x}+3}\ dx=[\frac{1}{\sqrt{3}}\tan^{-1}(\frac{e^x}{\sqrt{3}})]_0^\infty\]

OpenStudy (anonymous):

thats right solved it be letting e^x=u?

OpenStudy (anonymous):

u = e^x du = e^x dx:

OpenStudy (anonymous):

Yep, and this is arctangent basically: \(\large\frac{1}{u^2+\sqrt{3}}\)

OpenStudy (anonymous):

But ugh... the evaluation...

OpenStudy (anonymous):

sorry what is ugh?

OpenStudy (anonymous):

:-P It's an expression of disgruntlement

OpenStudy (fwizbang):

arctan(infinity)= pi/2.

OpenStudy (anonymous):

Grr is more angry frustrated, ugh is more fatigued frustrated

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

But that's e\(^{\infty}\)

OpenStudy (fwizbang):

One infinity is much like another here....

OpenStudy (anonymous):

Order of operations @fwizbang , agreed?

OpenStudy (anonymous):

Oh so you're saying it doesn't matter, point.

OpenStudy (anonymous):

*calculating*

OpenStudy (fwizbang):

You only need top worry about "different" infinities when you're dividing things.

OpenStudy (anonymous):

\(\large\frac{\pi}{3\sqrt3}\)?

OpenStudy (fwizbang):

why the 3 in the denom?(outside the root?)

OpenStudy (anonymous):

Greatest common denominator right?

OpenStudy (anonymous):

(27)^(1/2) = 9^(1/2) * (3)^(1/2)

OpenStudy (fwizbang):

\[\tan^{-1} (e^x/\sqrt{3}) -> \pi/2\] the root 3 makes no difference.

OpenStudy (anonymous):

Ty for the help gentlemen! :-)